OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. -20, -16, -12, -8, ...

ganeshie8 (ganeshie8):

its an 'arithmetic sequence', so lets find the common-difference first :)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

plus 4

ganeshie8 (ganeshie8):

Yes !

ganeshie8 (ganeshie8):

common-difference \(d = 4\)

ganeshie8 (ganeshie8):

see that, first term \(a = -20\)

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

simply plug those in the 'nth term of arithmetic sequence formula' :) do u have the formula in ur notes ?

ganeshie8 (ganeshie8):

here is the formula :- \(\large a_n = a+ (n-1)d\)

ganeshie8 (ganeshie8):

plugin a & d values we found before

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

an=20+(n-1)4

ganeshie8 (ganeshie8):

\(\large a_n=\color{red}{-}20+(n-1) \times 4\)

OpenStudy (anonymous):

an = -20 + 4(n - 1)

ganeshie8 (ganeshie8):

u forgot the negative sign before 20, otherwise it looks good :)

ganeshie8 (ganeshie8):

looks good :)

OpenStudy (anonymous):

or an = -20 x 4(n - 1)

ganeshie8 (ganeshie8):

its just an = -20 + 4(n - 1)

OpenStudy (anonymous):

thank you!!

ganeshie8 (ganeshie8):

np :)